Green Hab Report – Jan 12th

Crew 202 Greenhab Report 12-Jan-2019
Greenhab Officer:
Jake Qiu & Hermes Bolivar

Environmental Control:
Heating
Cooling w/ ambient air (1hrs)

40% Shade Cloth on
80% Shade Cloth on

Average Temperatures:
Low: 24.6°C
High: 39.5°C

Hours of Supplemental Light:
5

Daily Water Usage of Crops:
12 gallons

Water in Blue Tank (gallons):
51%, 151.5 gallons

Times of Watering for Crops:
1044
2031

Changes to crops:
* Threw away 2 pots of spinach due to it dying
* Radishes growing very well and one look almost ready for harvest
* Tomatoes looking lackluster
* Cucumbers still look bad, may need to move these pots to different locations
* Other crops display same progress

Narrative:
Watered plants throughout the day as needed. Showing Hermes from crew 203 the greenhouse for handover.
Experiment Results:
Date: SOL14 Crew 202 (SOL14 overall)

Watered Crew 201 experiments

Harvest:

Support/Supplies Needed:

Best Regards,

Jake

Sol Summary Report – Jan 12th

Crew 202 Sol Summary Report 12-Jan-2019

Sol: 14

Summary Title: Back on Earth

Author’s name: Cesare Guariniello and Oscar Ojeda

Mission Status: Nominal, the crew is back on Earth and met crew 203!

Sol Activity Summary: After breakfast, the crew began cleaning the habitat and the other facilities and got ready to meet crew 203. We had a short excursion with the Program Director and after dinner crew 203 is shadowing crew 202 for report training.

After a long trip, crew 203 was finally reunited in Grand Junction, had an amazing trip all the way to the station, and were warmly welcomed by crew 202. We joined the excursion and had our last dinner on Earth as we prepare for our mission

Look Ahead Plan: Tomorrow crew 202 will leave campus and crew 203 will receive field training

Anomalies in work: None

Weather: Sunny and warmer

Crew Physical Status: In great health

EVA: none

Reports to be filed: Sol Summary, Operations Report, Greenhab Report, Astronomy Report, Journalist Report

Support Requested: None

Operations Report – Jan 12th

Crew 202 Operations Report 12-JAN-2019

SOL: 14

Name of person filing report: Kasey Hilton

Non-nominal systems: Suit #9

Notes on non-nominal systems: One of suit #9’s electrical connections was loose yesterday and had to be soldered, but after soldering the suit would not turn on

Generator (hours run): 17hr 15min; Turned on last night (11Jan2019) at 17:15; Turned off this morning (12Jan2019) at 10:30; Turned on tonight (12Jan2019) at 17:29

Solar SOC – Turned on (11Jan2019) 81%; Turned off (12Jan2019) 100%; Turned on (12Jan2019) 83%

Diesel Reading – 60%

Propane Reading – 31%

Ethanol Free Gasoline – Not in use

Water (auxiliary tank) – Not in use

Water (static tank) – 65%; 345 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not applicable

Water in GreenHab – About 50%; 155 gallons

Water (loft) – At level marker 12

Static to Loft Pump used – Yes; At 20:19 to refill tank

Water Meter: 01400068

Toilet tank emptied: Yes, still very very stinky ://

Deimos rover used: No, still not functional

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: No

Hours: 67.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Opportunity rover used: No, brakes are still not working correctly

Hours: 45.6

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: No

Hours: 69.0

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Brake fluid in Opportunity was replaced but the brakes are still not functioning properly

ATV’s Used: None (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: None

Oil Added? No

ATV Fuel Used: None

# Hours the ATVs were Used today: None

Notes on ATVs: Nothing to report

HabCar used and why, where? Used to get water and dinner in town

CrewCar used and why, where? Used to pick up crew 203 in Grand Junction and get dinner in town

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Suit #9’s electrical connection in the recharger was soldered last night but did not turn on again after, even though the battery has sufficient charge; Suit #7 has not been fixed yet; The new crew engineer has been made aware of the situation

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

One last YEEHAW,

Kasey